OpenStudy (anonymous):

This graph of a function is a translation of y=2/x. What is a equation of the function?

OpenStudy (anonymous):

a. y=2/(x-6) - 3 b. y=2/(x-6) + 3 c. y=2/(x-3) - 6 d. y=2/(x-3) + 6